OpenStudy (anonymous):

If one x-intercept for the parabola y=4x^2-4x-8 is (2,0), what is the other..?

OpenStudy (anonymous):

do you know how to solve the equation?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

ok first simplify by dividing by 4 4(x^2 - x - 2) = 0 x^x - 2x + x - 2 = 0 x(x - 2) + 1(x-2) = 0 (x - 2)(x +1) = 0 x = 2 or -1 the other x-intercept is at x = -1 answer is (-1,0)

OpenStudy (anonymous):

there u go

OpenStudy (anonymous):

Thanks :) helps a lot